I would like some help on this please:
The formula for a regression equation is Y’ = 2X + 9.
a. What would be the predicted score for a person scoring 6 on X?
Does that mean 2 * 6 + 9 = 21 so Y'=21?
b. If someone’s predicted score was 14, what was this person’s score on X?
in that case it would be 2*14 + 9 = 37 so Y' = 37?
I'm hoping it's that easy :)

Answer by ewatrrr(24785)

regression equation is Y’ = 2X + 9
a. x = 6
y' = 2*6 + 9 = 21
|
b. y' = 14
14 = 2x + 9
5/2 = x = 2.5 0r 3 as a whole number